# Marine Saliency Segmenter: Object-Focused Conditional Diffusion with Region-Level Semantic Knowledge Distillation

### v1

@@ Abstract
Marine Saliency Segmentation (MSS) plays a pivotal role in various vision-based marine exploration tasks. However, existing marine segmentation techniques face the dilemma of object mislocalization and imprecise boundaries due to the complex underwater environment. Meanwhile, despite the impressive performance of diffusion models in visual segmentation, there remains potential to further leverage contextual semantics to enhance feature learning of region-level salient objects, thereby improving segmentation outcomes. Building on this insight, we propose DiffMSS, a novel marine saliency segmenter based on the diffusion model, which utilizes semantic knowledge distillation to guide the segmentation of marine salient objects. Specifically, we design a region-word similarity matching mechanism to identify salient terms at the word level from the text descriptions. These high-level semantic features guide the conditional feature learning network in generating salient and accurate diffusion conditions with semantic knowledge distillation. To further refine the segmentation of fine-grained structures in unique marine organisms, we develop the dedicated consensus deterministic sampling to suppress overconfident missegmentations. Comprehensive experiments demonstrate the superior performance of DiffMSS over state-of-the-art methods in both quantitative and qualitative evaluations.
@@ 1 Introduction
Marine Saliency Segmentation (MSS) focuses on segmenting visually salient objects within complex underwater environments to meet the growing requirement for fine-grained object recognition [12, 97, 38, 3]. Functionally, accurate recognition of marine instances contributes to applications like organism identification [10, 37], autonomous navigation [49, 100], and object detection [19, 4, 5]. However, the raw images captured directly by underwater vehicles tend to lose visual saliency, presenting various types of degradation, such as color distortion, low contrast, and blurred details [11, 22, 30, 96, 76]. Underwater degraded images with these defects usually exhibit indistinguishable object boundaries and a camouflaged appearance.
@@
With advances in large-scale annotated datasets [18, 89] and deep network architectures, many saliency detection methods [83, 35, 64, 99, 70, 71, 87] in natural image domains have made remarkable performances. However, they face challenges in underwater environments, where the poor visibility and fine-grained structures of marine organisms (e.g., fish, corals) greatly degrade the accuracy [98]. Existing MSS methods [14, 25, 27, 20, 42] follow the basic paradigm of a learning-based backbone and a decoder for segmentation. However, they usually stack multiple convolutional sequences with limited representational power into deep networks to extract deep features that require a lot of computational resources [23, 19]. Without well-designed backbones, they remain vulnerable to visual degradation and suffer from inaccurate boundary segmentation.
@@
Given the specific challenges posed by the MSS task, we explore the diffusion model [61] as a fitting solution due to its strong generative capabilities with well-defined conditional inputs. Despite the impressive performance of diffusion models with conditional prompts in visual segmentation tasks [59, 55, 7, 77], the potential of leveraging contextual semantics to generate diffusion conditions remains underexplored. While the contextual semantics usually come from the caption of the image, word-level concepts within the text can provide more direct fine-grained information about the salient object [74, 95]. In fact, by extracting and aligning these word-level semantic and visual features, we can guide the diffusion model to focus on these key salient regions and improve its segmentation accuracy.
@@
Motivated by the aforementioned analysis, we propose DiffMSS, an innovative diffusion-based marine saliency segmenter that leverages a region-level knowledge distillation scheme to guide the segmentation of marine salient objects. As depicted in Fig. 2, we first design a Word-level Semantic Saliency Extraction (WSSE) to adaptively identify salient terms described in the given text through region-word similarity matching. Then, these high-level text features transfer contextual semantic information to the Conditional Feature Learning Network (CFLN) based on Semantic Knowledge Distillation (SKD), guiding it to generate region-level visual features as object-focused diffusion conditions. This distillation scheme enables DiffMSS to improve inference efficiency without relying on WSSE and SKD for semantic saliency transfer during the testing phase. To further refine the segmentation of fine-grained structures, we develop a dedicated Consensus Deterministic Sampling (CDS) to suppress inaccurate segmentation caused by overconfidence in camouflaged marine instances.
@@ 2.1 Marine Saliency Segmenatation
Existing MSS methods can be roughly divided into handcrafted feature-based methods [28, 56, 19, 29, 86, 32] and deep learning-based methods [31, 27, 25, 23, 79, 80, 38]. Early handcrafted feature-based methods relied on low-level visual features to achieve segmentation [9, 53]. With the rise and advancement of visual foundation models, various network architectures have been proposed to address MSS. Li et al. [38] proposed a feature interaction encoder and cascaded decoder to extract more comprehensive information, while Liu et al. [44] combined channel and spatial attention modules to refine feature maps to obtain better object boundaries. Although these CNN-based models are effective, they cannot capture the long-range dependencies of complex marine objects and ignore the connectivity between discrete pixels [45]. Recently, instead of linearly stacking multiple convolutional layers in the network, several deep learning-based USD methods [2, 23, 43, 21, 14] incorporated visual transformers with wider receptive fields into their deep architectures. This way alleviates the computational burden brought by convolution to some extent, but these methods are unreliable in capturing saliency information by improving the encoder architecture.
@@ 2.2 Text-supervised Feature Matching
With the rise of text-supervised semantic segmentation, many studies [82, 74, 40, 78, 57, 81, 1] have utilized text prompts to enhance segmentation performance. These large vision-language models [58, 39, 41, 63] have been trained on large text-image datasets such as LAION-5B [60], enabling them to understand the alignment between text descriptions and visual elements. They train image and text encoders to align image-text pairs within a joint embedding space, thereby generating segmentation results with zero-shot supervision [52, 13, 94, 93]. Although straightforward, images may contain multiple object instances, and the semantic features of the text should match corresponding segments rather than the entire image. Several region-text alignment methods [34, 62, 1] have been proposed to strengthen the consistency between the segmented region and the text description, which enables the network to focus on segmenting the relevant regions described in the text.
@@ 2.3 Diffusion-based Image Segmentation
With their powerful generative capabilities, diffusion models have achieved impressive performance in terms of image restoration [90, 91, 65], object detection [6, 72, 92, 73], and depth estimation [84, 51, 33]. By leveraging the adaptive characteristics of the diffusion process, diffusion models have shown potential in various segmentation tasks [66, 24, 75, 47, 55]. For instance, DiffuMask [75] utilizes cross-modal attention maps between image features and conditional text embeddings to segment the most prominent object indicated by text prompts. LD-ZNet [55] performs text-based synthetic image segmentation by revealing rich semantic information within its internal features. However, existing diffusion models employ pixel-level corruption to generate the noised mask directly from the GT, which causes the model to mistakenly assume that the restored contours from the noised mask are accurate [67]. In addition, these methods often produce conditional features with limited discriminative representation. To address this, we propose a conditional feature learning network under the guidance of region-level semantic knowledge distillation to robustly generate discriminative conditional features.
@@ 3.2 Region-Level Semantic Saliency Knowledge Distillation
$\mathcal{L}_{\textit{consist}}=1-\frac{{Proj}(\mathcal{F}_{l}^{v})_{v}\cdot{Proj}(\mathcal{F}^{st})_{t}}{\|{Proj}(\mathcal{F}_{l}^{v})_{v}\|\,\|{Proj}(\mathcal{F}^{st})_{t}\|},$ (6)
@@ 4.1 Experimental Setups
$F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $m_{IoU}\uparrow$ $M_{AE}$downarrow $F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $m_{IoU}\uparrow$ $M_{AE}$$\downarrow$ $F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $m_{IoU}\uparrow$ $M_{AE}$downarrow
SUIM-Net [25] 0.783 0.856 0.797 0.628 0.1011 0.807 0.867 0.826 0.705 0.0787 0.711 0.733 0.679 0.639 0.1326
RMFormer [14] 0.828 0.910 0.867 0.746 0.0439 0.830 0.908 0.859 0.753 0.0623 0.790 0.851 0.796 0.711 0.1085
BBRF [48] 0.799 0.875 0.853 0.731 0.0462 0.816 0.891 0.856 0.748 0.0679 0.739 0.746 0.759 0.683 0.1119
TC-USOD [23] 0.907 0.968 0.922 0.872 0.0201 0.879 0.951 0.893 0.827 0.0388 0.856 0.917 0.859 0.816 0.0631
RobustSAM [8] 0.872 0.947 0.908 0.861 0.0396 0.859 0.918 0.865 0.806 0.0573 0.839 0.893 0.847 0.802 0.0717
DualSAM [88] 0.917 0.968 0.924 0.881 0.0185 0.876 0.937 0.881 0.817 0.0465 0.858 0.921 0.861 0.813 0.0637
DiffMSS (Ours) 0.923 0.967 0.932 0.893 0.0189 0.891 0.947 0.908 0.835 0.0376 0.867 0.927 0.873 0.821 0.0566
@@
Evaluation Metrics. We adopt five commonly used metrics for MSS tasks evaluation, including weighted F-measure ($F_{\beta}^{w}$) [50], max E-measure ($E_{\phi}^{m}$) [17], S-measure ($S_{\alpha}$) [16], Mean Intersection over Union ($m_{IoU}$) [46], and mean absolute error ($M_{AE}$) [54].
@@ 4.2 Comparison with State-of-the-Arts
[…]
SUIM-Net [25] 12.22 71.46 0.265 0.1035
RMFormer [14] 174.19 563.14 0.315 0.0598
BBRF [48] 74.01 31.13 0.107 0.0629
TC-USOD [23] 117.64 29.64 0.089 0.0318
RobustSAM [8] 407.76 1492.60 0.137 0.0490
DualSAM [88] 159.95 318.05 0.081 0.0322
DiffMSS (Ours) 68.41 24.98 0.033 0.0296
[…]
@@ 4.3 Evaluation of Model Efficiency
Figure 7: Visual ablation of segmentation results with (-w/ SKD) and without (-w/o SKD) Semantic Knowledge Distillation.
Figure 8: Conditional features ablations of with (-w/ SKD) and without (-w/o SKD) Semantic Knowledge Distillation at different sampling steps.
@@
$\checkmark$ 0.860 0.932 0.905 0.1110
$\checkmark$ 0.923 0.967 0.932 0.0189
Table 3: Ablation study of Semantic Knowledge Distillation.
@@
$\checkmark$ 0.923 0.967 0.932 0.0189
4.4 Ablation Study
Ablation Study of Semantic Knowledge Distillation. We first conduct a visual ablation study with and without semantic knowledge distillation (denoted as “-w/ SKD” and “-w/o SKD”). Fig. 7 shows that the absence of caption guidance in “-w/o SKD” leads to over-segmentation or under-segmentation, while Fig. 8 presents that “-w/ SKD” contributes to generating fine-grained conditional features. In the quantitative evaluation in Table 3, “-w SKD” outperforms “-w/o SKD” in four indicators, which indicates the necessity of SKD for salient object segmentation.
@@ 4.4 Ablation Study
$\checkmark$ 0.664 0.853 0.796 0.0719 65.23 19.49
$\checkmark$ $\checkmark$ 0.831 0.943 0.899 0.0399 66.62 20.31
$\checkmark$ $\checkmark$ $\checkmark$ 0.883 0.955 0.908 0.0229 67.59 21.66
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.923 0.967 0.932 0.0189 68.41 24.98
@@
$\checkmark$ $\checkmark$ 0.860 0.932 0.895 0.1109
$\checkmark$ $\checkmark$ 0.892 0.951 0.919 0.0247
$\checkmark$ $\checkmark$ 0.856 0.925 0.873 0.1369
$\checkmark$ $\checkmark$ $\checkmark$ 0.923 0.967 0.932 0.0189
Table 6: Ablation study of loss function terms.
@@
$\checkmark$ 0.886 0.935 0.906 0.0274
$\checkmark$ 0.923 0.967 0.932 0.0189
Table 7: Ablation study of Consensus Deterministic Sampling.

### v2

@@ Abstract
Marine Saliency Segmentation (MSS) plays a pivotal role in various vision-based marine exploration tasks. However, existing techniques often face the dilemma of imprecise boundaries due to the interference-rich nature of underwater environments, where suspended particles, low contrast, and color distortion hinder accurate segmentation. Meanwhile, despite the impressive performance of diffusion models in visual tasks, there remains potential to further leverage contextual semantics to enhance feature learning of region-level salient objects, thereby improving segmentation outcomes. Building on this insight, we propose DiffMSS, a novel marine saliency segmenter based on the diffusion model, which utilizes semantic knowledge distillation to guide the detection of marine salient objects. Specifically, we design a region-word similarity matching mechanism to identify salient terms at the word level from the text descriptions. These high-level semantic features guide the conditional feature learning network in generating salient and accurate diffusion conditions with semantic knowledge distillation. To further refine the segmentation of fine-grained structures in unique marine organisms, we develop a dedicated consensus deterministic sampling to suppress overconfident missegmentations. Extensive experiments demonstrate the superior performance of DiffMSS over state-of-the-art methods in both quantitative and qualitative evaluations.
@@ 1 Introduction
Marine Saliency Segmentation (MSS) focuses on segmenting visually salient objects within complex underwater environments to meet the growing requirement for fine-grained object recognition [60]. Functionally, accurate recognition of marine instances contributes to applications like organism identification [24], autonomous navigation [32], and object detection [3]. However, the raw images captured directly by underwater vehicles tend to lose visual saliency, presenting various types of degradation, such as color distortion, low contrast, and blurred details [59]. Underwater degraded images with these defects usually exhibit indistinguishable object boundaries and a camouflaged appearance.
@@
With advances in large-scale annotated datasets [12, 57] and deep network architectures, many saliency detection methods [53, 23, 44] in natural image domains have made remarkable performances. However, they face challenges in underwater environments, where the poor visibility and fine-grained structures of marine organisms (e.g., fish, corals) greatly degrade the accuracy [61]. Existing MSS methods [8, 17, 19, 14, 27] follow the basic paradigm of a learning-based backbone and a decoder for segmentation. However, they usually stack multiple convolutional sequences with limited representational power into deep networks to extract deep features that require a lot of computational resources [16, 13]. Without well-designed backbones, they remain vulnerable to visual degradation and suffer from inaccurate boundary segmentation.
@@
Given the specific challenges posed by the MSS task, we explore the diffusion model [42] as a fitting solution due to its strong generative capabilities. Despite the impressive performance of diffusion models [37, 4] with conditional prompts in common segmentation tasks, the potential of leveraging contextual semantics to generate the diffusion conditions remains underexplored. Moreover, these models usually adopt coarse-grained image-level or text-level features as conditions for diffusion (as shown in Fig. 1). While image-level captions provide contextual information, word-level concepts can offer more precise semantic cues related to salient regions [50]. By extracting and aligning these word-level semantics with visual features, we can guide the diffusion model to focus on these key salient regions and improve its segmentation accuracy.
@@
Motivated by the aforementioned analysis, we propose DiffMSS, an innovative diffusion-based marine saliency segmenter that leverages a region-level knowledge distillation scheme to guide the detection of marine salient objects. As depicted in Fig. 2, we first design a word-level semantic saliency extraction to adaptively identify salient terms described in the given text through region-word similarity matching. Then, these high-level text features transfer contextual semantic information to the conditional feature learning network based on semantic knowledge distillation, guiding it to generate region-level visual features as object-focused diffusion conditions. To refine the segmentation of fine-grained structures, we develop a dedicated consensus deterministic sampling to suppress inaccurate segmentation caused by overconfidence in camouflaged marine objects.
@@ 2.1 Marine Saliency Segmenatation
Existing MSS methods can be roughly divided into handcrafted feature-based methods [20, 38] and deep learning-based methods [21, 19, 17, 16, 25]. Early handcrafted feature-based methods relied on low-level visual features to achieve segmentation [6, 35]. With the rise and advancement of visual foundation models, various network architectures have been proposed to address MSS. Li et al. [25] proposed a feature interaction encoder and cascaded decoder to extract more comprehensive information, while Liu et al. [29] combined channel and spatial attention modules to refine feature maps to obtain better object boundaries. Although these CNN-based models are effective, they cannot capture the long-range dependencies of complex marine objects and ignore the connectivity between discrete pixels [30]. Recently, instead of linearly stacking multiple convolutional layers in the network, several deep learning-based USD methods [2, 16, 28, 15, 8] incorporated visual transformers with wider receptive fields into their deep architectures. This way alleviates the computational burden brought by convolution to some extent, but these methods are unreliable in capturing saliency information by improving the encoder architecture.
@@ 2.2 Text-supervised Feature Matching
With the rise of text-supervised semantic segmentation, many studies [52, 50, 39, 1] have utilized text prompts to enhance segmentation performance. These large vision-language models [40, 26] have been trained on large text-image datasets such as LAION-5B [41], enabling them to understand the alignment between text descriptions and visual elements. They train image and text encoders to align image-text pairs within a joint embedding space, thereby generating segmentation results with zero-shot supervision [34, 7]. Although straightforward, images may contain multiple object instances, and the semantic features of the text should match corresponding segments rather than the entire image. Several region-text alignment methods [22, 43] have been proposed to strengthen the consistency between the segmented region and the text description, which enables the network to focus on segmenting the relevant regions described in the text.
@@ 2.3 Diffusion-based Image Segmentation
With their powerful generative capabilities, diffusion models have achieved impressive performance in terms of image restoration [45], object detection [58], and depth estimation [54]. By leveraging the adaptive characteristics of the diffusion process, diffusion models have shown potential in various segmentation tasks [46, 37]. For instance, DiffuMask [51] utilizes cross-modal attention maps between image features and conditional text embeddings to segment the most prominent object indicated by text prompts. LD-ZNet [37] performs text-based synthetic image segmentation by revealing rich semantic information within its internal features. However, existing diffusion models employ pixel-level corruption to generate the noised mask directly from the GT, which causes the model to mistakenly assume that the restored contours from the noised mask are accurate [47]. In addition, these methods often produce conditional features with limited discriminative representation. To address this, we propose a conditional feature learning network under the guidance of region-level semantic knowledge distillation to robustly generate discriminative conditional features.
@@ 3.2 Region-Level Semantic Saliency Knowledge Distillation
$\mathcal{L}_{\textit{consist}}=-\frac{1}{N}\sum_{i=1}^{N}\frac{{Proj}(\mathcal{F}_{l}^{v}(i))_{v}\cdot{Proj}(\mathcal{F}^{st}(i))_{t}}{\|{Proj}(\mathcal{F}_{l}^{v}(i))_{v}\|_{2}\,\|{Proj}(\mathcal{F}^{st}(i))_{t}\|_{2}},$ (6)
@@ 4.1 Experimental Setups
$F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $M_{AE}$downarrow $F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $M_{AE}$$\downarrow$ $F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $M_{AE}$downarrow
SUIM-Net [17] 0.783 0.856 0.797 0.1011 0.807 0.867 0.826 0.0787 0.734 0.751 0.739 0.1162
RMFormer [8] 0.828 0.910 0.867 0.0439 0.830 0.908 0.859 0.0623 0.829 0.865 0.817 0.0942
BBRF [31] 0.902 0.935 0.913 0.0317 0.856 0.891 0.856 0.0679 0.847 0.876 0.839 0.0695
TC-USOD [16] 0.910 0.953 0.912 0.0236 0.879 0.951 0.893 0.0388 0.856 0.917 0.859 0.0631
RobustSAM [5] 0.897 0.946 0.909 0.0356 0.861 0.924 0.869 0.0567 0.839 0.893 0.847 0.0717
DualSAM [56] 0.909 0.959 0.916 0.0218 0.876 0.937 0.881 0.0465 0.858 0.921 0.861 0.0637
DiffMSS 0.912 0.956 0.922 0.0203 0.891 0.947 0.908 0.0376 0.867 0.927 0.873 0.0566
@@
Evaluation Metrics. We adopt five commonly used metrics for MSS tasks evaluation, including weighted F-measure ($F_{\beta}^{w}$) [33], max E-measure ($E_{\phi}^{m}$) [11], S-measure ($S_{\alpha}$) [10], and mean absolute error ($M_{AE}$) [36].
@@ 4.3 Evaluation of Model Efficiency
[…]
SUIM-Net [17] 12.22 71.46 0.265 0.1006
RMFormer [8] 174.19 563.14 0.315 0.0503
BBRF [31] 74.01 31.13 0.107 0.0385
TC-USOD [16] 117.64 29.64 0.089 0.0287
RobustSAM [5] 407.76 1492.60 0.214 0.0409
DualSAM [56] 159.95 325.68 0.088 0.0280
DiffMSS (Ours) 68.41 24.98 0.033 0.0253
4.4 Ablation Study
Ablation Study of Semantic Knowledge Distillation. We conduct an ablation study with and without semantic knowledge distillation (“-w/ SKD” and “-w/o SKD”) and evaluate segmentation performance using image-text matching (“I-T”) or region-word matching (“R-W”) schemes in the “-w/ SKD” case. Table 3 shows that utilizing “I-T” matching can improve the performance of saliency segmentation to a certain extent, but our “R-W” matching scheme achieves the highest scores across all metrics, which demonstrates the effectiveness of word-level semantic alignment in achieving object-focused diffusion.
Table 3: Ablation study of Semantic Knowledge Distillation.
@@ 4.4 Ablation Study
$\checkmark$ 0.897 0.942 0.913 0.0355
$\checkmark$ 0.912 0.956 0.922 0.0203
@@
$\checkmark$ 0.912 0.956 0.922 0.0203
@@
$\checkmark$ 0.664 0.853 0.796 0.1219 65.23 19.49
$\checkmark$ $\checkmark$ 0.831 0.943 0.879 0.0868 66.62 20.31
$\checkmark$ $\checkmark$ $\checkmark$ 0.893 0.945 0.910 0.0292 67.59 21.66
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.912 0.956 0.922 0.0203 68.41 24.98
@@
Table 6: Ablation study of loss function terms.
@@
$\checkmark$ $\checkmark$ 0.897 0.942 0.903 0.0355
$\checkmark$ $\checkmark$ 0.885 0.941 0.899 0.0378
$\checkmark$ $\checkmark$ 0.856 0.925 0.873 0.0706
$\checkmark$ $\checkmark$ $\checkmark$ 0.912 0.956 0.922 0.0203
@@
Table 7: Ablation study of Consensus Deterministic Sampling.
@@
$\checkmark$ 0.903 0.938 0.916 0.0267
$\checkmark$ 0.912 0.956 0.922 0.0203